Solve:
m cubed plus 3m squared minus 54x =0

Answer by richwmiller(17219)

m^3+3m^2-54m=0
Are you sure it isn't -54m?
m (m^2+3 m-54) = 0
(m-6) m (m+9) = 0
m=6 and m=-9 and m=0